# A New Upper Bound for Distributed Hypothesis Testing Using the Auxiliary Receiver Approach

### v3

@@ III-A Strict improvement for Gaussian Sources
where $\rho_{0}\in[-1,1]$, and without loss of generality, $\rho_{1}\in[0,1]$. Even though we stated our bounds for discrete settings, a similar argument goes through for jointly Gaussian random variables. Moreover, in this section, we measure the communication rate in nats per symbol and use the natural logarithm.
III-A1 Existing Bounds
@@
$\rho\triangleq\left\{\begin{array}[]{l l}\frac{\rho_{0}-\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{1}\cup\mathcal{D}_{2}$}\\ \frac{\rho_{0}+\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in$\mathcal{D}_{3}$}.\\ \end{array}\right.$
$C\triangleq\left\{\begin{array}[]{l l}0&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{1}\cup\mathcal{D}_{2}$}\\ \frac{2\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{3}$}.\\ \end{array}\right.$
@@
$E_{SHA}(R)=\sup_{\eta\geq 0}\max\Big{[}E_{\eta}(R),\max_{0\leq\beta\leq\beta_{\max}}$
$\min\{E_{\eta}(\beta R),\overline{E}_{\eta}(\beta R)-(\beta-1)R+E_{\eta}(\beta R){\mathbf{1}}_{\{\rho_{0}<\rho_{1}\}}\}\Big{]}$
@@
$E_{\eta}(R)=D\Big{(}\frac{(\rho_{1}-\eta)^{2}\sigma_{R}^{2}+\rho_{1}^{2}(1-\sigma_{R}^{2})}{1+\eta(\eta-2\rho_{1})\sigma_{R}^{2}},\frac{1+\eta(\eta-2\rho_{0})\sigma_{R}^{2}}{1+\eta(\eta-2\rho_{1})\sigma_{R}^{2}}\Big{)}$
$\overline{E}_{\eta}(R)=D\Big{(}\frac{\eta^{2}\sigma_{R}^{2}}{1+\eta^{2}\sigma_{R}^{2}},\frac{1+\eta(\eta-2\rho_{0})\sigma_{R}^{2}}{1+\eta^{2}\sigma_{R}^{2}}\Big{)}$
$\sigma_{R}^{2}=1-\exp(-2R)$
$D(\gamma,P)=\frac{\gamma+P-A(\gamma,P)}{2(1-\gamma)}-\frac{1}{2}\log\frac{A(\gamma,P)-1+\gamma}{2\gamma}$
$A(\gamma,P)=\sqrt{(1-\gamma)^{2}+4P\gamma}$
@@ III-A2 New Bound
$\mathcal{R}$ =$=\big{\{}(P_{Z|XY},Q_{Z|XY}):Q_{YZ|X}=Q_{Z|X}Q_{Y|Z},$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad P_{XZ}=Q_{XZ}\big{\}}$
@@
$D(P_{YZ}\|Q_{YZ})$ $=\frac{g^{2}-{\rho_{1}}{g}[(\rho_{0}-bg)g+b]}{{g^{2}}-{\rho_{1}^{2}}}-1$
$\qquad\qquad+\frac{1}{2}\log\frac{{g^{2}}-{\rho_{1}^{2}}}{{g^{2}}-{g^{2}}[(\rho_{0}-bg)g+b]^{2}},$
@@
$g\in[-1,1],\quad b^{2}(1-g^{2})\leq 1-\rho_{0}^{2},\quad((\rho_{0}-bg)g+b)^{2}\leq 1,$
$\rho_{1}^{2}\leq g^{2}.$
Plots comparing the bound with all the other known bounds are given in Fig. 2 and Fig. 3. One can observe that our bound is always better than the bound in [16] for these parameters. Computing the bound in [20] is difficult, so we did not replicate the simulation done by [20] and just extracted the points from the curve given in [20]. One can observe that the auxiliary receiver bound is better than the bound in [20] when $\rho_{0}$ is not close to one. For $\rho_{0}$ close to one, the bound in [20] is better, which suggests multiple auxiliary receivers can yield strictly better bounds for large $\rho_{0}$. We have discussed combining the idea in [20] with our technique in Section III-B.
@@
$I_{P}(Y,U|Z)$
$=\frac{1}{2}\log\left\{\frac{1-(ag+b)^{2}}{1-(a+bg)^{2}(1-\sigma^{2})-\frac{[(a+bg)g\sigma^{2}+b(1-g^{2})]^{2}}{1-(1-\sigma^{2})g^{2}}}\right\},$
@@
We show the optimality of a jointly Gaussian distribution on $(U,X)$ when evaluating $G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$ (and hence the above weaker upper bound) in Appendix A.
If the variance of $P_{X|U}$ is $\sigma^{2}\leq 1$. The derivative of the denominator of $I_{P}(Y,U|Z)$ is $\frac{a^{2}(g^{2}-1)^{2}}{(g^{2}\sigma^{2}-g^{2}+1)^{2}}\geq 0$. The optimal $\sigma^{2}$ that maximizes $I_{P}(Y,U|Z)$ within feasible region $\mathcal{T}_{R}$ is then given by:
@@
III-B The bound in [20] and a sequence of auxiliary receivers
@@
$=\log\frac{\sigma_{Q,Y}}{\sigma_{P,Y}}+\frac{\sigma_{P,Y}^{2}-\sigma^{2}_{Q,Y}+k^{2}}{2\sigma^{2}_{Q,Y}}$
$-\log\frac{\sigma_{Q,Z}}{\sigma_{P,Z}}-\frac{\sigma_{P,Z}^{2}-\sigma^{2}_{Q,Z}+k^{2}}{2\sigma^{2}_{Q,Z}}$
@@
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$ (22)
$+E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$ (23)
@@
$\leq G_{P_{XY},Q_{XY},P_{XY^{\prime}},Q_{XY^{\prime}}}(R)$
$+G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)+E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$
@@
Appendix A Evaluation of the upper bound in the Gaussian setting
@@
$G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$
$=\sup_{P_{U|X}}\mathbb{E}_{P_{U}}\left[\max_{\hat{Q}_{X}}D(P_{Y|U}\|\hat{Q}_{Y})-D(P_{Z|U}\|\hat{Q}_{Z})\right]$
where the supremum is over $P_{U|X}$ satisfying $I_{P}(U;X|Y)\leq R$ and $I_{P}(U;X|Z)\leq R$. Assume that $X$ is a scalar standard Gaussian random variable under both hypotheses. Assume that $P_{Y|X},Q_{Y|X},P_{Z|X}$ and $Q_{Z|X}$ are additive Gaussian noise channels with additive noise variances $\sigma_{P,Y|X}^{2}$, $\sigma_{Q,Y|X}^{2}$, $\sigma_{P,Z|X}^{2}$, $\sigma_{Q,Z|X}^{2}$ respectively. If $\sigma_{Q,Z|X}>\sigma_{Q,Y|X}$, the above expression becomes infinity. This can be seen by choosing a constant random variable $U$ and following the choice in Example 2 given in Section III-B.
@@
$=-h_{P}(Y|U)+h_{P}(Z|U)+\mathbb{E}_{P_{U}}\max_{\hat{Q}_{X}}\mathbb{E}_{P_{YZ|U}}\log\frac{\hat{Q}_{Z}}{\hat{Q}_{Y}}$
@@
Let $G$ be a standard Gaussian random variable, independent of $(X,Y,Z)$, and (under either hypothesis) let
@@
$(U^{\prime},X^{\prime},\tilde{X}^{\prime},\tilde{Y}^{\prime},\tilde{Z}^{\prime})$
$:=((U^{*}_{1},U^{*}_{2}),(X^{*}_{1},X^{*}_{2}),(\tilde{X}^{*}_{1},\tilde{X}^{*}_{2}),(\tilde{Y}^{*}_{1},\tilde{Y}^{*}_{2}),(\tilde{Z}^{*}_{1},\tilde{Z}^{*}_{2}))$
@@
$\begin{pmatrix}X_{+}\\ X_{-}\end{pmatrix}:=\begin{pmatrix}\sqrt{t}&\sqrt{1-t}\\ -\sqrt{1-t}&\sqrt{t}\end{pmatrix}\begin{pmatrix}X^{*}_{1}\\ X^{*}_{2}\end{pmatrix}$
@@
$=\mathbb{E}_{P_{U^{\prime}}}\Big{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{+}\tilde{Y}_{-}|U^{\prime}}\|\hat{Q}_{Y_{+}Y_{-}})$ (31)
$\quad-D(P_{\tilde{Z}_{+}\tilde{Z}_{-}|U^{\prime}}\|\hat{Q}_{Z_{+}Z_{-}})\Big{]}-\delta h_{P}(\tilde{Y}_{+}\tilde{Y}_{-}|U^{\prime})$
@@
$\leq\mathbb{E}_{P_{U^{\prime}}}\bigg{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{-}|U^{\prime}\tilde{Y}_{+}}\|\hat{Q}_{\tilde{Y}_{-}|\tilde{Y}_{+}}|P_{\tilde{Y}_{+}|U^{\prime}})$ (32)
$\quad-D(P_{\tilde{Z}_{-}|U^{\prime}\tilde{Y}_{+}}\|\hat{Q}_{\tilde{Z}_{-}|\tilde{Y}_{+}}|P_{\tilde{Y}_{+}|U^{\prime}})\bigg{]}$
$\quad+\mathbb{E}_{P_{U^{\prime}}}\bigg{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{+}|U^{\prime}\tilde{Z}_{-}}\|\hat{Q}_{\tilde{Y}_{+}|\tilde{Z}_{-}}|P_{\tilde{Z}_{-}|U^{\prime}})$ (33)
$\quad-D(P_{\tilde{Z}_{+}|\tilde{Z}_{-}U^{\prime}}\|\hat{Q}_{\tilde{Z}_{+}|\tilde{Z}_{-}}|P_{\tilde{Z}_{-}|U^{\prime}})\bigg{]}$
@@
$+\mathbb{E}_{P_{\hat{U}^{\prime}_{+}}}\bigg{[}\max_{\hat{Q}_{X_{+}}}D(P_{\tilde{Y}_{+}|\hat{U}^{\prime}_{+}}\|\hat{Q}_{\tilde{Y}_{+}})-D(P_{\tilde{Z}_{+}|\hat{U}^{\prime}_{+}}\|\hat{Q}_{\tilde{Z}_{+}})\bigg{]}$
$\qquad-\delta h_{P}(\tilde{Y}_{+}|U^{\prime}_{+})-\delta h_{P}(\tilde{Y}_{-}|U^{\prime}_{-})-\delta I_{P}(\tilde{Y}_{+};\tilde{Z}_{-}|U^{\prime}).$ (34)

### v4

@@ III-C Strict improvement for Gaussian Sources
where $\rho_{0}\in[-1,1]$, and without loss of generality, $\rho_{1}\in[0,1]$. Even though we stated our bounds for discrete settings, a similar argument goes through for jointly Gaussian random variables. Moreover, in this section, we measure the communication rate in nats per symbol, the logarithms and the exponent are in base $e$.
III-C1 Existing Bounds
@@
$\rho\triangleq\left\{\begin{array}[]{l l}\frac{\rho_{0}-\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{1}\cup\mathcal{D}_{2}$},\\ \frac{\rho_{0}+\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{3}$}.\\ \end{array}\right.$
$C\triangleq\left\{\begin{array}[]{l l}0&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{1}\cup\mathcal{D}_{2}$},\\ \frac{2\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{3}$}.\\ \end{array}\right.$
@@
$E_{SHA}(R)=\sup_{\eta\geq 0}\max\Big{[}E_{\eta}(R),\max_{0\leq\beta\leq\beta_{\max}}\min\{E_{\eta}(\beta R),\overline{E}_{\eta}(\beta R)-(\beta-1)R+E_{\eta}(\beta R){\mathbf{1}}_{\{\rho_{0}<\rho_{1}\}}\}\Big{]}$
@@
$E_{\eta}(R)=D\Big{(}\frac{(\rho_{1}-\eta)^{2}\sigma_{R}^{2}+\rho_{1}^{2}(1-\sigma_{R}^{2})}{1+\eta(\eta-2\rho_{1})\sigma_{R}^{2}},\frac{1+\eta(\eta-2\rho_{0})\sigma_{R}^{2}}{1+\eta(\eta-2\rho_{1})\sigma_{R}^{2}}\Big{)},$
$\overline{E}_{\eta}(R)=D\Big{(}\frac{\eta^{2}\sigma_{R}^{2}}{1+\eta^{2}\sigma_{R}^{2}},\frac{1+\eta(\eta-2\rho_{0})\sigma_{R}^{2}}{1+\eta^{2}\sigma_{R}^{2}}\Big{)},$
$\sigma_{R}^{2}=1-\exp(-2R),$
$D(\gamma,P)=\frac{\gamma+P-A(\gamma,P)}{2(1-\gamma)}-\frac{1}{2}\log\frac{A(\gamma,P)-1+\gamma}{2\gamma},$
$A(\gamma,P)=\sqrt{(1-\gamma)^{2}+4P\gamma}.$
@@ III-C2 New Bound
$\mathcal{R}$ =$=\big{\{}(P_{Z|XY},Q_{Z|XY}):Q_{YZ|X}=Q_{Z|X}Q_{Y|Z},P_{XZ}=Q_{XZ}\big{\}},$
@@
$D(P_{YZ}\|Q_{YZ})$ $=\frac{g^{2}-{\rho_{1}}{g}[(\rho_{0}-bg)g+b]}{{g^{2}}-{\rho_{1}^{2}}}-1+\frac{1}{2}\log\frac{{g^{2}}-{\rho_{1}^{2}}}{{g^{2}}-{g^{2}}[(\rho_{0}-bg)g+b]^{2}},$
@@
$g\in[-1,1],\quad b^{2}(1-g^{2})\leq 1-\rho_{0}^{2},\quad((\rho_{0}-bg)g+b)^{2}\leq 1,\quad\rho_{1}^{2}\leq g^{2}.$
Plots comparing the bound with all the other known bounds are given in Fig. 6 and Fig. 7. One can observe that our bound is always better than the bound in [9] for these parameters. Computing the bound in [11] is difficult, so we did not replicate the simulation done by [11] and just extracted the points from the curve given in [11]. One can observe that the auxiliary receiver bound is better than the bound in [11] when $\rho_{0}$ is not close to one. For $\rho_{0}$ close to one, the bound in [11] is better, which suggests multiple auxiliary receivers can yield strictly better bounds for large $\rho_{0}$. We have discussed combining the idea in [11] with our technique in Section III-D.
@@
$I_{P}(Y;U|Z)=\frac{1}{2}\log\left\{\frac{1-(ag+b)^{2}}{1-(a+bg)^{2}(1-\sigma^{2})-\frac{[(a+bg)g\sigma^{2}+b(1-g^{2})]^{2}}{1-(1-\sigma^{2})g^{2}}}\right\},$
@@
We show the optimality of a jointly Gaussian distribution on $(U,X)$ when evaluating $G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$ (and hence the above weaker upper bound) in Appendix B.
If the variance of $P_{X|U}$ is $\sigma^{2}\leq 1$. The derivative of the denominator of $I_{P}(Y;U|Z)$ is $\frac{a^{2}(g^{2}-1)^{2}}{(g^{2}\sigma^{2}-g^{2}+1)^{2}}\geq 0$. The optimal $\sigma^{2}$ that maximizes $I_{P}(Y;U|Z)$ within feasible region $\mathcal{T}_{R}$ is then given by:
@@
III-D The bound in [11] and a sequence of auxiliary receivers
@@
$=\log\frac{\sigma_{Q,Y}}{\sigma_{P,Y}}+\frac{\sigma_{P,Y}^{2}-\sigma^{2}_{Q,Y}+k^{2}}{2\sigma^{2}_{Q,Y}}-\log\frac{\sigma_{Q,Z}}{\sigma_{P,Z}}-\frac{\sigma_{P,Z}^{2}-\sigma^{2}_{Q,Z}+k^{2}}{2\sigma^{2}_{Q,Z}}.$
@@
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)+E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R).$ (40)
@@
$\leq G_{P_{XY},Q_{XY},P_{XY^{\prime}},Q_{XY^{\prime}}}(R)+G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)+E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$
@@
Appendix A Alternative characterization of the upper bound of Rahman and Wagner
The following alternative characterization of the Rahman-Wagner bound for discrete alphabet random variables is useful in our numerical evaluations:
Lemma 3 (Alternative characterization of the bound in Theorem 3).
For discrete sources, the bound in Theorem 3 can be equivalently written as
$\min_{(P_{Z|XY},Q_{Z|XY})\in\mathcal{R}(P_{XY},Q_{XY})}\left\{D(P_{YZ}\|Q_{YZ})+\min_{\lambda\in[0,1]}\max_{P_{U|X}}[I_{P}(Y;U|Z)+\lambda(R-I_{P}(X;U|Z))]\right\}.$
where the maximum is over all $P_{U|X}$.
Proof of Lemma 3.
The Rahman-Wagner bound can be expressed as
$\min_{(P_{Z|XY},Q_{Z|XY})\in\mathcal{R}(P_{XY},Q_{XY})}\left\{D(P_{YZ}\|Q_{YZ})+\max_{P_{U|X}}\min_{\lambda\geq 0}[I_{P}(Y;U|Z)+\lambda(R-I_{P}(X;U|Z))]\right\}.$
To prove the lemma, it suffices to show the following two equalities:
$\max_{P_{U|X}}\min_{\lambda\geq 0}[I_{P}(Y;U|Z)+\lambda(R-I_{P}(X;U|Z))]$ $=\min_{\lambda\geq 0}\max_{P_{U|X}}[I_{P}(Y;U|Z)+\lambda(R-I_{P}(X;U|Z))]$ (41)
$=\min_{\lambda\in[0,1]}\max_{P_{U|X}}[I_{P}(Y;U|Z)+\lambda(R-I_{P}(X;U|Z))].$ (42)
We first show the equality in (42). For any $\lambda\geq 1$,
$I(Y;U|Z)-I(X;U|Z)+\lambda R-(\lambda-1)I(X;U|Z)$
$=-I(X;U|Y,Z)+\lambda R-(\lambda-1)I(X;U|Z)$
$\leq\lambda R.$
Moreover, $\lambda R$ is achieved when $U$ is a constant random variable. Therefore, for any $\lambda\geq 1$, we have
$\max_{P_{U|X}}[I_{P}(Y;U|Z)+\lambda(R-I_{P}(X;U|Z))]=\lambda R$
and is minimized when $\lambda=1$.
It remains to show the equality in (41). Let
$\mathcal{A}=\{(a_{1},a_{2}):a_{1}=I(X;U|Z),a_{2}=I(Y;U|Z),\quad\text{for some }p(u|x)\}.$
The equality in (41) can be expressed as
$\max_{(a_{1},a_{2})\in\mathcal{A}}\min_{\lambda\geq 0}[a_{2}+\lambda(R-a_{1})]$ $=\min_{\lambda\geq 0}\max_{(a_{1},a_{2})\in\mathcal{A}}[a_{2}+\lambda(R-a_{1})].$ (43)
We use Sion’s Minimax Theorem [39] to show that we can exchange the order of $\max$ and $\min$. We can use Sion’s minimax theorem because
  • The function $a_{2}+\lambda(R-a_{1})$ is linear in $a_{1}$, $a_{2}$, and $\lambda$.
  • The domain set $\{\lambda:\lambda\geq 0\}$ is a convex set.
  • The set $\mathcal{A}$ is a closed, compact and convex set. To see this, observe that the set $\mathcal{A}$ is compact because $a_{1}\in[0,H(X|Z)]$ and $a_{2}\in[0,H(Y|Z)]$; it is closed because one can impose a cardinality bound of $|\mathcal{U}|\leq|\mathcal{X}|+2$ on the auxiliary random variable $U$ using the standard arguments based on the Caratheodery’s theorem. The set $\mathcal{A}$ is convex because for any $p_{U_{1}|X}$ and $p_{U_{2}|X}$, we can take $Q\in\{1,2\}$ to be a Bernouli (1/2) random variable that is independent of all previously defined variables and define $U=(Q,U_{Q})$.
Appendix B Evaluation of the upper bound in the Gaussian setting
@@
$G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)=\sup_{P_{U|X}}\mathbb{E}_{P_{U}}\left[\max_{\hat{Q}_{X}}D(P_{Y|U}\|\hat{Q}_{Y})-D(P_{Z|U}\|\hat{Q}_{Z})\right]$
where the supremum is over $P_{U|X}$ satisfying $I_{P}(U;X|Y)\leq R$ and $I_{P}(U;X|Z)\leq R$. Assume that $X$ is a scalar standard Gaussian random variable under both hypotheses. Assume that $P_{Y|X},Q_{Y|X},P_{Z|X}$ and $Q_{Z|X}$ are additive Gaussian noise channels with additive noise variances $\sigma_{P,Y|X}^{2}$, $\sigma_{Q,Y|X}^{2}$, $\sigma_{P,Z|X}^{2}$, $\sigma_{Q,Z|X}^{2}$ respectively. If $\sigma_{Q,Z|X}>\sigma_{Q,Y|X}$, the above expression becomes infinity. This can be seen by choosing a constant random variable $U$ and following the choice in Example 2 given in Section III-D.
@@
$=-h_{P}(Y|U)+h_{P}(Z|U)+\mathbb{E}_{P_{U}}\max_{\hat{Q}_{X}}\mathbb{E}_{P_{YZ|U}}\log\frac{\hat{Q}_{Z}}{\hat{Q}_{Y}},$
@@
[…]
@@
$(U^{\prime},X^{\prime},\tilde{X}^{\prime},\tilde{Y}^{\prime},\tilde{Z}^{\prime})=((U^{*}_{1},U^{*}_{2}),(X^{*}_{1},X^{*}_{2}),(\tilde{X}^{*}_{1},\tilde{X}^{*}_{2}),(\tilde{Y}^{*}_{1},\tilde{Y}^{*}_{2}),(\tilde{Z}^{*}_{1},\tilde{Z}^{*}_{2}))$
@@
$\begin{pmatrix}X_{+}\\ X_{-}\end{pmatrix}=\begin{pmatrix}\sqrt{t}&\sqrt{1-t}\\ -\sqrt{1-t}&\sqrt{t}\end{pmatrix}\begin{pmatrix}X^{*}_{1}\\ X^{*}_{2}\end{pmatrix}$
@@
$=\mathbb{E}_{P_{U^{\prime}}}\Big{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{+}\tilde{Y}_{-}|U^{\prime}}\|\hat{Q}_{Y_{+}Y_{-}})-D(P_{\tilde{Z}_{+}\tilde{Z}_{-}|U^{\prime}}\|\hat{Q}_{Z_{+}Z_{-}})\Big{]}-\delta h_{P}(\tilde{Y}_{+}\tilde{Y}_{-}|U^{\prime})$
@@
$\leq\mathbb{E}_{P_{U^{\prime}}}\bigg{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{-}|U^{\prime}\tilde{Y}_{+}}\|\hat{Q}_{\tilde{Y}_{-}|\tilde{Y}_{+}}|P_{\tilde{Y}_{+}|U^{\prime}})-D(P_{\tilde{Z}_{-}|U^{\prime}\tilde{Y}_{+}}\|\hat{Q}_{\tilde{Z}_{-}|\tilde{Y}_{+}}|P_{\tilde{Y}_{+}|U^{\prime}})\bigg{]}$
$\quad+\mathbb{E}_{P_{U^{\prime}}}\bigg{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{+}|U^{\prime}\tilde{Z}_{-}}\|\hat{Q}_{\tilde{Y}_{+}|\tilde{Z}_{-}}|P_{\tilde{Z}_{-}|U^{\prime}})-D(P_{\tilde{Z}_{+}|\tilde{Z}_{-}U^{\prime}}\|\hat{Q}_{\tilde{Z}_{+}|\tilde{Z}_{-}}|P_{\tilde{Z}_{-}|U^{\prime}})\bigg{]}$
@@
$\quad+\mathbb{E}_{P_{\hat{U}^{\prime}_{+}}}\bigg{[}\max_{\hat{Q}_{X_{+}}}D(P_{\tilde{Y}_{+}|\hat{U}^{\prime}_{+}}\|\hat{Q}_{\tilde{Y}_{+}})-D(P_{\tilde{Z}_{+}|\hat{U}^{\prime}_{+}}\|\hat{Q}_{\tilde{Z}_{+}})\bigg{]}$
$\quad-\delta h_{P}(\tilde{Y}_{+}|U^{\prime}_{+})-\delta h_{P}(\tilde{Y}_{-}|U^{\prime}_{-})-\delta I_{P}(\tilde{Y}_{+};\tilde{Z}_{-}|U^{\prime}).$ (51)
@@
Acknowledgements
This work was supported by the CUHK Direct Grant 4055193. The authors thank Chandra Nair for inspiring discussions and for helping with the proof of the optimality of Gaussian distributions.